# Learning to combine top-down context and feed-forward representations under ambiguity with apical and basal dendrites

### v1

@@ 4.1 Ambiguous dataset generation
To create the dataset for our task, we first train a conditional variational autoencoder (CVAE) on standard MNIST/EMNIST [64, 98]. The purpose of the CVAE in this setup is to use the learned latent space to generate images that are ambiguous between two digits. While category-free interpolation in the latent space can be done with a vanilla VAE, the CVAE approach facilitates the generation of ambiguous images conditioned on the specified categories. We then concatenate the input to the decoder (latent $z$) with a linear combination of one-hot vectors representing the target digits. We define an image to be ambiguous when it lies at the $50\%\pm 5\%$ decision boundary between two classes of a classifier trained on standard MNIST/EMNIST. We collect a set of images in triplets, where two images are unambiguous from unique classes $y_{0},y_{1}$, and the third is ambiguous (using our criterion) between $y_{0},y_{1}$, generated by the CVAE. This dataset format simplifies the implementation of our task setup.
4.2 Model architecture
The model consists of three components: the pretrained backbone, the readout for classification, and the topdown network. We first train a variational autoencoder with a convolutional encoder and decoder to form smooth latent representations of unambiguous characters. Once the variational autoencoder backbone is trained, its weights are frozen and a readout (MLP) is trained to classify the latent representations.
@@ 4.2.1 Oracle context
In the oracle case, top-down signals are provided in the form of a one-hot vector:
@@ 4.2.2 Temporal context
In the temporal case, context signals were provided by a Gated Recurrent Unit (GRU) network [60]. We trained a GRU with hidden state of size 128 to predict the modulo 10 sum of MNIST digits in a sequence of two digits’. The input to the GRU at time $t$ is not $x_{t}$ , but rather $\mathbf{Ub}_{t}$, encoded by the pretrained backbone. After processing a sequence of digits ($x_{t-2},x_{t-1}$), the final output state $o_{t}$ of the GRU was then used as a top-down context signal.
@@ 4.3.1 Loss function
The topdown network $g(x,c|\Theta_{a})$ is explicitly trained to minimize a loss function composed of a sum over the following 5 scenarios: ambiguous input with matching context (ambig.Match), unambiguous input with matching context (unambig.Match), unambiguous input with irrelevant context (unambig.Irrel), ambiguous input with irrelevant context (ambig.Irrel), and unambiguous input with contradictory context (unambig.Contra). All scenarios are therefore optimized together for every gradient update step. We use mean squared error as the loss criterion between the predicted latents ( computed as in equation 2) and the target latents for each scenario. Each minibatch is sampled from dataset $D$, and scenarios sampled from the minibatch and set of contexts $C$.
@@
It should be noted that $x_{0},x_{1}$ are unambiguous images sampled from $D$ The optimum of $\Theta_{a}$ is given by:
@@
$\mu_{s}^{*}:=\begin{cases}\mathbf{U}\mathbf{\sigma(b_{c_{match}})},&\text{if $s\in\{1\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})},&\text{if $s\in\{2,3,5\}$}\\ \mathbf{U}\mathbf{\sigma(b_{ambig})},&\text{if $s\in\{4\}$}\\ \end{cases}$ (7)
Where $b_{c}=f(x_{c}|\Theta_{b})$, and $b_{unambig}=f([x_{0},x_{1}]|\Theta_{b})$. Predicted latents are given by:
@@
Where $\mathbf{a_{b,c}}={g(\mathbf{Ub,c}|\Theta_{a})}$ (See pseudocode 1)
@@
4:for triplet $\texttt{triplet minibatch}(\mathbf{x_{0}},\mathbf{y_{0}}),\mathbf{x_{ambig}},(\mathbf{x_{1}},\mathbf{y_{1}})\texttt{ sampled from }D$ do
5:     $y\leftarrow\{\mathbf{y_{0}},\mathbf{y_{1}}\}$
6:     $x\leftarrow\{\mathbf{x_{0},x_{1}}\}$
7:     # Concatenate all scenarios into single batch
8:     $\mathbf{c_{match}}\leftarrow random\_choice(y,size=B)$ Sample context uniformly from y over minibatch
@@
10:     $\mathbf{c_{contra}}\leftarrow\mathbf{y_{1}}$
11:     # combine all vectors into one for efficiency
12:     $c\leftarrow cat(c_{match},c_{match},c_{irrel},c_{irrel},c_{contra})$
13:     $\mu^{*}\leftarrow cat(\mu_{c_{match}},\mu_{y},\mu_{y},\mu_{ambig},\mu_{y_{0}})$
14:     $x\leftarrow(x_{ambig},x,x,x_{ambig},x_{0})$
15:     $b\leftarrow f(x;\Theta_{b})$
16:     $a\leftarrow g(\mathbf{U}b\oplus c;\Theta_{a})$
17:     $\mathbf{\hat{h}}\leftarrow\textbf{relu}(b)\odot(1+\textbf{relu}(a))$
18:     $\mathbf{\hat{\mu}}\leftarrow\mathbf{U\hat{h}}$
19:     $Loss\leftarrow\textbf{mse}(\mathbf{\hat{\mu}},\mu^{*})$
20:     # autograd optimizer step
21:     $\Theta_{a}\leftarrow\Theta_{a}+\lambda\nabla_{\Theta_{a}}Loss$

### v2

@@ 4.1 Ambiguous dataset generation
We developed a new visual contextual integration task composed of multiple scenarios that incorporate varying levels of ambiguity of sensory input and informativeness of contextual information.
To create the image dataset for our task, we used a digit-conditioned generative approach based on conditional variational autoencoders (CVAEs). First, we train a CVAE on standard a image dataset, either MNIST for digits or EMNIST for characters (Doersch, 2021; Sohn et al., 2015). We then use the learned latent space along with a digit-conditioned input vector to generate images that are ambiguous between two digits by interpolating between the one-hot vectors of the two digits.
We then assess the ambiguous samples using a trained classifier on the original dataset (either MNIST or EMNIST), and discard images that are outside the $50\%(\pm 5\%)$ decision boundary between the two digits. We organize our dataset of images into triplets, where two images are unambiguous from digits $y_{0},y_{1}$, and the third is generated by the CVAE to be ambiguous between $y_{0},y_{1}$. This dataset format simplifies the implementation of our task setup.
4.2 Contextual integration model architecture
The contextual integration model consists of three components: the pretrained backbone that computes representations of input images, the readout network for classifying their respective representations, and the topdown network which modulates the intermediate representations from the pretrained backbone based on contextual information. The backbone is implemented as a convolutional variational autoencoder (VAE), and was pretrained on unambiguous characters with the standard ELBO loss. VAEs are known to learn a smooth latent space, which we assume here to be an important ingredient to allow learning of top-down contextual modulations. Next, we freeze the VAE weights and train a readout (MLP) on the VAEs latent representations of images with a classification objective (cross entropy loss). Finally, we freeze these components and train the topdown network in two contextual tasks, with oracle and temporal context respectively, as described in the next sections. The topdown network $g(\cdot)$ is implemented as a multilayer perceptron (MLP) mapping the concatenated VAE latents and context vector to apical activations (eq. 3), which then modulate the VAEs intermediate representation. Various modulation functions (hadamard, additive, concatenation) are explored in the main text.
@@ 4.2.1 Oracle context
In the oracle case, contextsignals are provided to the top-down network $g$ in the form of a one-hot vector:
@@ 4.2.2 Temporal context
In the temporal case, the information is presented in a sequence, and preceding elements determine the contextual information available to the model at the current time $t$. Therefore, we leverage a Gated Recurrent Unit (GRU) network (Cho et al., 2014) to dynamically capture and represent the temporal dependencies in the sequence. We trained our GRU to predict the modulo 10 sum of MNIST digits in a sequence of two digits’, reflecting the task in our sequential MNIST/EMNIST dataset. An important detail is that the input to the GRU at time $t$ is not the image $x_{t}$ , but rather the latent $\mathbf{Ub}_{t}$, encoded by the pretrained backbone. After processing a sequence of digits ($x_{t-2},x_{t-1}$), the output state $o_{t}$ of the GRU was then used as a top-down context signal, in place of the one-hot context from the oracle case. We used a hidden state of size 128 for the GRU.
@@ 4.3.1 Loss function
The topdown network $g(x,c|\Theta_{a})$ is explicitly trained to minimize a loss function jointly across the following 5 conditions: ambiguous input with matching context (ambig.Match), unambiguous input with matching context (unambig.Match), unambiguous input with irrelevant context (unambig.Irrel), ambiguous input with irrelevant context (ambig.Irrel), and unambiguous input with contradictory context (unambig.Contra). We use mean squared error as the loss criterion between the predicted (computed as in equation 2) and the target latent representations. Each minibatch is sampled from dataset $D$, and training data for each condition are equally balanced in the minibatch.
@@
$x_{0},x_{1}$ are unambiguous images sampled from $D$.
The optimum of $\Theta_{a}$ is given by:
@@
$\mu_{s}^{*}:=\begin{cases}\mathbf{U}\mathbf{\sigma(b_{c_{match}})},&\text{if $s\in\{1\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})},&\text{if $s\in\{2,3\}$}\\ \mathbf{U}\mathbf{\sigma(b_{ambig})},&\text{if $s\in\{4\}$}\\ \mathbf{U}\mathbf{\sigma(b_{c_{match}})},&\text{if $s\in\{5\}$}\end{cases}$ (7)
Where $b_{c}=f(x_{c}|\Theta_{b})$ and
$b_{unambig}=f([x_{0},x_{1}]|\Theta_{b})$.
Predicted latents are given by:
@@
Where $\mathbf{a_{b,c}}={g(\mathbf{Ub,c}|\Theta_{a})}$.
@@
3:for triplet $\texttt{triplet minibatch}(\mathbf{x_{0}},\mathbf{y_{0}}),\mathbf{x_{ambig}},(\mathbf{x_{1}},\mathbf{y_{1}})\texttt{ sampled from }D$ do
4:     $y\leftarrow cat(\mathbf{y_{0}},\mathbf{y_{1}})$
5:     $x\leftarrow cat(\mathbf{x_{0},x_{ambig},x_{1}})$
6:     $\mathbf{c_{match}},\mathbf{c_{contra}}\leftarrow random\_choice(y,size=(2,B))\texttt{ \# Sample context from y without replacement}$
@@
8:     $c\leftarrow cat(c_{match},c_{irrel},c_{contra})$
[…]
10:     $\mu=\mathbf{U}b$
11:     $\mu c\leftarrow\mu\oplus c$
12:     $a\leftarrow g(\mu c;\Theta_{a})$
[…]
15:     # autograd optimizer step
16:     $\Theta_{a}\leftarrow\Theta_{a}+\lambda\nabla_{\Theta_{a}}Loss$